<e> Solve the first order linear inhomogeneous differential equation using the bernoulli method

(2x+1)y,=4x+2y



1
Expert's answer
2022-02-25T10:30:29-0500

"(2x+1)y'=4x+2y\\
y=u(x)v(x)\\
(2x+1)(u'v+uv')=4x+2uv\\
(2x+1)u'v+(2x+1)uv'-2uv=4x\\
(2x+1)u'v+u((2x+1)v'-2v)=4x\\
1) (2x+1)v'-2v=0\\
\frac{dv}{v}=\frac{2dx}{2x+1}\\
\int\frac{dv}{v}=\int\frac{2dx}{2x+1}\\
\ln|v|=\ln|2x+1|\\
v=2x+1\\
2) (2x+1)u'\cdot (2x+1)=4x\\
du=\frac{4x}{(2x+1)^2}\\
u=\int\frac{4x}{(2x+1)^2}dx=\\
=\int\frac{2(2x+1-1)}{(2x+1)^2}dx=\\
=\int(\frac{2}{2x+1}-\frac{2}{(2x+1)^2})dx=\\
=\ln|2x+1|+\frac{1}{2x+1}+C\\
y= (2x+1)(\ln|2x+1|+\frac{1}{2x+1}+C)"